# Parallel methods for quasinonexpansive mappings in a Hilbert space

### v1

@@ 2. Preliminaries
[…]
@@
for all $x,y\in C$; $T$ is said to be strongly quasinonexpansive [pNACA2017, pNACA2015, MR4136436, MR3908455, MR3213161] if $T$ is quasinonexpansive and $Tx_{n}-x_{n}\to 0$ whenever $\{x_{n}\}$ is a bounded sequence in $C$ and $\left\lVert x_{n}-p\right\rVert-\left\lVert Tx_{n}-p\right\rVert\to 0$ for some point $p\in\operatorname{F}(T)$; $T$ is said to be demiclosed at 00 if $z\in C$ and $Tz=0$ whenever $\{x_{n}\}$ is a sequence in $C$ such that $x_{n}\rightharpoonup z$ and $Tx_{n}\to 0$; see, for example, [MR1074005]. It is clear that $I-T$ is demiclosed at 00 if and only if $\hat{\mathrm{F}}(T)=\operatorname{F}(T)$. Moreover, under the assumption that $C$ is closed and convex, we know the following:
@@
If a strongly nonexpansive mapping in the sense of [MR0470761, MR2581778, MR2799767, MR2671943, MR2377867] has a fixed point, then it is strongly quasinonexpansive above; see also [MR2581778, MR2799767, MR2671943, MR2377867].
@@
Let $T\colon C\to H$ be a mapping with a fixed point. Since $\beta_{n}-\alpha_{n}\to 0\Leftrightarrow\beta_{n}^{2}-\alpha_{n}^{2}\to 0$ for all bounded sequences $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ in $[0,\infty)$, $T$ is strongly quasinonexpansive if and only if $T$ is of type (sr) in the sense of [MR3258665, MR2529497, MR2884574]; see also [MR2358984, MR3745220].
@@
  • If $\inf_{n}\beta_{n}>0$, then $\{S_{n}\}$ is a strongly quasinonexpansive sequence;
@@
Let $H$ be a Hilbert space, $C$ a nonempty subset of $H$, $\{U_{n}\}$ be a sequence of mappings of $C$ into $H$, $N$ a positive integer, and $T_{i}$ a mapping of $C$ into $H$ for $i\in\Lambda$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Suppose that $\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$ is nonempty, and that for any $x\in C$ and $n\in\mathbb{N}$ there exists $k\in\arg\max\{\left\lVert T_{i}x-x\right\rVert\colon i\in\Lambda\}$ such that $U_{n}x=T_{k}x$. Then the following hold:
@@ 3. Parallel methods for quasinonexpansive mappings
Let $H$ be a Hilbert space, $C$ a nonempty closed convex subset of $H$, $N$ a positive integer, and $T_{i}\colon C\to H$ a quasinonexpansive mapping for $i\in\Lambda$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Let $F$ be the set of common fixed points of $\{T_{i}\}_{i\in\Lambda}$, that is, $F=\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
@@
for $x\in C$ and $n\in\mathbb{N}$, where $j(x)=\min M(x)$ for $x\in C$. Then we can verify that for any $x\in C$ and $n\in\mathbb{N}$, there exists $k\in M(x)$ such that $U_{n}x=T_{k}x$. Thus Lemma 2.11 implies that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, $\{U_{n}\}$ is a strongly quasinonexpansive sequence, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. Set $S_{n}=\beta_{n}I+(1-\beta_{n})U_{n}$ for $n\in\mathbb{N}$. Then Lemma 2.10 shows that $\{S_{n}\}$ is a strongly quasinonexpansive sequence and $\hat{\mathrm{F}}(\{S_{n}\})=\bigcap_{n}\operatorname{F}(S_{n})=F$. From the definition of $U_{n}$, it is clear that $x_{n+1}=P_{C}\bigl{(}\alpha_{n}u+(1-\alpha_{n})S_{n}x_{n}\bigr{)}$ for all $n\in\mathbb{N}$. Using Theorem 2.12, we conclude that $\{x_{n}\}$ converges strongly to $P_{F}(u)$. ∎
@@
Let $H$, $C$, $N$, $\Lambda$, $\{\alpha_{n}\}$, and $\{\beta_{n}\}$ be the same as in Theorem 3.1. Let $T_{i}\colon C\to C$ be a quasinonexpansive mapping for $i\in\Lambda$, $F$ the set of common fixed points of $\{T_{i}\}_{i\in\Lambda}$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
@@ 4. Parallel methods for strongly quasinonexpansive mappings
Let $H$, $N$, $S_{i}$, $\Lambda$, $F$ be the same as in Theorem 4.1 and $\{x_{n}\}$ a sequence defined by $x_{1}\in H$ and
@@
Let $M$ be a set-valued mapping of $H$ into $\Lambda$ defined by (4.1) for $x\in H$ and $U_{n}\colon H\to H$ a mapping defined by (4.2) for $x\in H$ and $n\in\mathbb{N}$. As in the proof of Theorem 4.1, it follows from Lemma 2.11 that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, $\{U_{n}\}$ is a strongly quasinonexpansive sequence, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. From the definition of $U_{n}$, it is clear that $x_{n+1}=U_{n}x_{n}$ for all $n\in\mathbb{N}$. By applying Theorem 2.13, the proof is complete. ∎

### v2

@@ 2. Preliminaries
Throughout the present paper, $H$ denotes a real Hilbert space, $\left\langle\,\cdot\,$$\cdot\,\right\rangle$ the inner product of $H$, $\left\lVert\,\cdot\,\right\rVert$ the norm of $H$, $C$ a nonempty subset of $H$, II the identity mapping on $H$, and $\mathbb{N}$ the set of positive integers. Strong convergence of a sequence $\{x_{n}\}$ in $H$ to $z\in H$ is denoted by $x_{n}\to z$ and weak convergence by $x_{n}\rightharpoonup z$.
@@
for all $x,y\in C$; $T$ is said to be strongly quasinonexpansive [pNACA2017, pNACA2015, MR4136436, MR3908455, MR3213161] if $T$ is quasinonexpansive and $Tx_{n}-x_{n}\to 0$ whenever $\{x_{n}\}$ is a bounded sequence in $C$ and $\left\lVert x_{n}-p\right\rVert-\left\lVert Tx_{n}-p\right\rVert\to 0$ for some point $p\in\operatorname{F}(T)$; $T$ is said to be demiclosed at 0 if $z\in C$ and $Tz=0$ whenever $\{x_{n}\}$ is a sequence in $C$ such that $x_{n}\rightharpoonup z$ and $Tx_{n}\to 0$; see, for example, [MR1074005]. It is clear that ITI-T is demiclosed at 0 if and only if $\hat{\mathrm{F}}(T)=\operatorname{F}(T)$. Moreover, under the assumption that $C$ is closed and convex, we know the following:
@@
If a strongly nonexpansive mapping in the sense of [MR0470761, MR2581778, MR2799767, MR2671943, MR2377867] has a fixed point, then it is strongly quasinonexpansive above.
@@
Let $T\colon C\to H$ be a mapping with a fixed point. Since $\beta_{n}-\alpha_{n}\to 0\Leftrightarrow\beta_{n}^{2}-\alpha_{n}^{2}\to 0$ for all bounded sequences $\{\alpha_{n}\}$ and $\{\beta_{n}\}$ in $0,\infty)$, $T$ is strongly quasinonexpansive if and only if $T$ is of type (sr) in the sense of [MR3258665, MR2529497, MR2884574]; see also [MR2358984, MR3745220].
@@
  • If each $U_{n}$ is quasinonexpansive and $\inf_{n}\beta_{n}>0$, then $\{S_{n}\}$ is a strongly quasinonexpansive sequence;
@@
Let $H$ be a Hilbert space, $C$ a nonempty subset of $H$, $\{U_{n}\}$ be a sequence of mappings of $C$ into $H$, NN a positive integer, and $T_{i}$ a mapping of $C$ into $H$ for $i\in\Lambda$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Suppose that $\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$ is nonempty, and that for any $x\in C$ and $n\in\mathbb{N}$ there exists $k\in\arg\max\{\left\lVert T_{i}x-x\right\rVert\colon i\in\Lambda\}$ such that $U_{n}x=T_{k}x$. Then the following hold:
@@ 3. Parallel methods for quasinonexpansive mappings
Let $H$ be a Hilbert space, $C$ a nonempty closed convex subset of $H$, NN a positive integer, and $T_{i}\colon C\to H$ a quasinonexpansive mapping for $i\in\Lambda$, where $\Lambda=\{i\in\mathbb{N}\colon 1\leq i\leq N\}$. Let $F$ be the set of common fixed points of $\{T_{i}\}_{i\in\Lambda}$, that is, $F=\bigcap_{i\in\Lambda}\operatorname{F}(T_{i})$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
@@
for $x\in C$ and $n\in\mathbb{N}$, where $j(x)=\min M(x)$ for $x\in C$. Then we can verify that for any $x\in C$ and $n\in\mathbb{N}$, there exists $k\in M(x)$ such that $U_{n}x=T_{k}x$. Thus Lemma 2.11 implies that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, each $U_{n}$ is quasinonexpansive, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. Set $S_{n}=\beta_{n}I+(1-\beta_{n})U_{n}$ for $n\in\mathbb{N}$. Then Lemma 2.10 shows that $\{S_{n}\}$ is a strongly quasinonexpansive sequence and $\hat{\mathrm{F}}(\{S_{n}\})=\bigcap_{n}\operatorname{F}(S_{n})=F$. From the definition of $U_{n}$, it is clear that $x_{n+1}=P_{C}\bigl{(}\alpha_{n}u+(1-\alpha_{n})S_{n}x_{n}\bigr{)}$ for all $n\in\mathbb{N}$. Using Theorem 2.12, we conclude that $\{x_{n}\}$ converges strongly to $P_{F}(u)$. ∎
@@
Let $H$, $C$, NN, $\Lambda$, $\{\alpha_{n}\}$, and $\{\beta_{n}\}$ be the same as in Theorem 3.1. Let $T_{i}\colon C\to C$ be a quasinonexpansive mapping for $i\in\Lambda$, $F$ the set of common fixed points of $\{T_{i}\}_{i\in\Lambda}$, and $\{x_{n}\}$ a sequence defined by $x_{1}\in C$ and
@@ 4. Parallel methods for strongly quasinonexpansive mappings
Let $H$, NN, $S_{i}$, $\Lambda$, $F$ be the same as in Theorem 4.1 and $\{x_{n}\}$ a sequence defined by $x_{1}\in H$ and
@@
Let $M$ be a set-valued mapping of $H$ into $\Lambda$ defined by (4.1) for $x\in H$ and $U_{n}\colon H\to H$ a mapping defined by (4.2) for $x\in H$ and $n\in\mathbb{N}$. As in the proof of Theorem 4.1, it follows from Lemma 2.11 that $F=\bigcap_{n}\operatorname{F}(U_{n})\neq\emptyset$, $\{U_{n}\}$ is a strongly quasinonexpansive sequence, and $\hat{\mathrm{F}}(\{U_{n}\})=\bigcap_{n}\operatorname{F}(U_{n})$. From the definition of $U_{n}$, it is clear that $x_{n+1}=U_{n}x_{n}$ for all $n\in\mathbb{N}$. Applying Theorem 2.13, we complete the proof. ∎